Some types of molecules that do not possess an intrinsic electric dipole moment can be given one by an external electric field in a process called charge separation, or polarization. In this process, their internal charge distribution becomes distorted by the field, which results in the region of a molecule on the side in the direction of the field gaining a positive net charge and the region on the other side gaining a negative net charge. Both charges have equal magnitudes, and the electric neutrality of the molecule as a whole is maintained. The electric field is said to induce an electric dipole moment in such a molecule. When the field is canceled, the molecule reverts to its unpolarized state and loses its electric dipole moment. The electric behavior of such a molecule can be modeled by a pair of ±1.60 x 10-19 C charges connected by a spring with force constant 0.000253 N/m. The spring must be imagined as possessing zero relaxed length so that normally the charges overlap and the molecule has zero dipole moment. In the presence of an electric field, however, the positive charge is pulled in the direction of the field, and the negative charge is pulled in the opposite direction. This charge separation endows the molecule with an electric dipole moment. Find the charge separation distance in an electric field of 5.07 × 105 N/C. charge separation distance:| 8.112 x10 What is the magnitude of the induced electric dipole moment? 32 electric dipole moment: 1.29792 ×10
Answer #1

The force on a charge from the dipole in the electric field can be calculated using

F = E= 1.6 x 10-19 x 5.07 x 105 = 8.112 x 10-14 N

At equilibrium, the tension in the spring balances this force. So, the spring force is

F = kr = 8.112 x 10-14 N

> C= 8.112 x 10-14 n -= 3.2 x 10-10 m 0.000253 N/m

So, the charge separation distance is 3.2 x 10-10 m.

The dipole moment is given by

p=qd = 1.6 x 10-19 X 3.2 x 10-10 = 5.12 x 10-29 C.m

So, the dipole moment is 5.12 x 10-29 C.m.